Mirrors CERN QMUL |
ATLAS Level 1 Calorimeter Trigger Software | |
L1Calo Software | Minutes: 6 April 2006 | |
![]() |
Software phone meeting on 6 April 2006Present: Bruce, Eric, Florian, Gilles, Kambiz, Murrough, Norman, Paolo, Richard, Steve, Tony, Victor, Weiming. Status reportsBirminghamGilles has added extra CPM status information in the IGUI. Meanwhile Chris has provided some new variable occupancy test vector patterns to help with full crate tests. After a change to the CMM module services, Gilles can now run CMM timing scans. A TTC problem has been observed where the broadcast for stop/start of the DSS seems to cause some clock instability. This seems to be a hardware problem and will be investigated further. Richard has been examining COOL extraction and browsing tools. He reported on two different tools. The first is COOL_IO which provides command line tools for extraction into ROOT or text files. However the available versions at CERN dont appear to have been updated for recent versions of COOL. The other tool is the KTIDBExplorer which is an update of the Lisbon groups conditions database browser with support for COOL. This is also available on lxplus though Richard had some problems installing it locally. He will circulate more details of both these tools to those interested. HeidelbergFlorian has changed the PPM VME readout buffer so it now has all the data from one event in one buffer. He has also added some PPM status information to the IGUI. Victor has committed ppmMonitoring code to analyse the cable test data. He is about to update ppmDecoder to adapt to Florians latest changes. Florian has circulated an email with some ideas about improving the handling of multistep runs. He would appreciate rapid feedback. MainzRainer reported that the JEM production, and hence testing, is delayed (as reported at the previous hardware phone meeting). QMULNothing new to report. RALNorman has been updating ROD documentation with recent programming model changes. He also mentioned that he and Richard had been investigating an problem observed with overwriting of the CMM parity mask registers. They will check if the method to update IS status is touching this register. Dave reported that he and Adam had discussed DCS with Jim Cook at CERN last week. They have agreed a scheme whereby the TCM can be made to look as much as possible like an ELMB. In some cases this requires packing/unpacking of multiple measurements into single CANopen longwords. StockholmNo report. CERNBruce said that SLC4 had now been certified and would soon become the default release at CERN. TDAQ will migrate at some point. Meanwhile Steve reported that all the simulation packages now work with the forthcoming TDAQ release. Database WorkshopThe dates for this are now fixed as Monday-Tuesday 22-23 May at CERN. Some people might like to stay on for a few days longer to make it a more general experts week. ROD and BUSYNorman reported that Robert Mclaren has asked ROD designers to provide a complete description of the use of ROD BUSY in each system. Norman will produce this, mainly an extract from the ROD spec, and circulate it before returning it to Robert. MonitoringPaolo has circulated a few documents from the monitoring working group (MWG) for us to comment on. For some of these we dont yet have enough experience to make detailed comments, but people are still encouraged to read them and give feedback via Paolo. One request from the MWG is to know how many monitoring PCs we expect to have in our racks downstairs. We have been assuming of the order of two or three such PCs. We certainly want something downstairs for commissioning before the monitoring farm upstairs is available. Comissioning DicussionNorman chaired a long discussion about the state of commissioning and calibration during which I failed to take notes. Next meetingsThe next phone meeting will be on Thursday 27 April at 11:00 CET (10:00 GMT). Last updated on 12-Apr-2006 by Murrough Landon |